Given an isosceles Triangle, whose one angle is 120 and inradius is √3. So area of triangle is?

Grade:11

1 Answers

Arun
6 years ago
Hello Student,
 
Area= root3 b^2 /4
=>sin120/a=sin30/b
=>a=root3 *b
Area=s*root3
s=(a+2b)/2=(a+2b)*root3*(1/2)
Area=12+7*root3
 
Thanks & Regards
Arun(askIITians forum expert)